given f(1/2)=2x^+3x-6 what is the answer
Answer by checkley77(12844)

f(1/2)=2x^2+3x-6 I'll assume there is a 2 after the 2x^.
f(1/2)=2(1/2)^2+3*1/2-6
f(1/2)=2*1/4+3/2-6
f(1/2)=1/2+3/2-6
f(1/2)=4/2-6
f(1/2)=2-6
f(1/2)=-4 ans.